OpenStudy (chris215):

If f and g are differentiable functions for all real values of x such that f(1) = 4, g(1) = 3, f '(3) = −5, f '(1) = −4, g '(1) = −3, g '(3) = 2, then find h '(1) if h(x) = f(x)/g(x)

OpenStudy (anonymous):

\[\left(\frac{f}{g}\right)'=\frac{gf'-fg'}{g^2}\] you have all the number you need, plug them in

OpenStudy (solomonzelman):

By the quotient rule: \(\large\color{black}{ \displaystyle \frac{d}{dx} \left[\frac{f(x)}{g(x)}\right] =\frac{f'(x)g(x)-f(x)g'(x)}{\left\{g(x)\right\}^2}}\) At x=1 (or h(1)), you get: \(\large\color{black}{ \displaystyle h'(1)=\frac{f'(1)g(1)-f(1)g'(1)}{\left\{g(1)\right\}^2}}\)

OpenStudy (solomonzelman):

then, you are given that: f(1)=4 g(1)=3 f '(1) = -4 g '(1) = -3

OpenStudy (solomonzelman):

Go for it...

OpenStudy (chris215):

so 0?

OpenStudy (solomonzelman):

Yes. f '(1) g(1) = f(1) g '(1) (-4) ( 3) = (4) ( -3) -12=-12

OpenStudy (solomonzelman):

-12-(-12)=-12+12=0

OpenStudy (solomonzelman):

very nice

OpenStudy (chris215):

thanks so much :)

OpenStudy (solomonzelman):

yw